# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v2

@@ 5.3.4. Request completion
4:vattention.init_cache(config_params)
@@
5.4. Optimizations
@@
There are two primary challenges in using CUDA virtual memory support for serving LLMs. First, cuMemCreate currently allocates a minimum of 2MB physical memory page. Large pages can waste physical memory due to internal fragmentation. Second, invoking CUDA APIs incurs high latency. This section details a set of simple-yet-effective optimizations that we introduce to overcome these limitations.
5.4.1. Mitigating internal fragmentation
We mitigate internal fragmentation by reducing the granularity of physical memory allocation. NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB (mcmgpus, ; gpu-mismanaged, ; gpu-reverse-engg-tlb1, ; pascalmmu, ). Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes. The simplest way to achieve this would be to extend the existing CUDA virtual memory APIs (listed in Table 2) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes). Unfortunately, the CUDA APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
Fortunately, some part of NVIDIA drivers (particularly related to unified memory management) is open-source. Therefore, we implement a new set of APIs in the open-source NVIDIA drivers to mimic the same functionality that existing CUDA APIs provide but with support for multiple page sizes. The second column in Table 2 shows our new APIs: most of our APIs have a one-to-one relationship with existing CUDA APIs except for vMemMap that combines the functionality of cuMemMap and cuMemSetAccess, and vMemRelease that combines the functionality of cuMemUnmap and cuMemRelease for simplicity. In contrast to CUDA APIs, our APIs can allocate memory in 64KB, 128KB and 256KB page sizes. A serving framework can configure a desired page size in vAttention while initializing it. The last set of columns in Table 2 shows the latency of each API with different page sizes.
5.4.2. Hiding memory allocation latency
@@
The serving framework invokes the step API in every iteration. The latency of step depends on how many new pages need to be mapped in the virtual tensors of KV-cache. Consider, for example, that the KV-cache of one request needs to be extended for Yi-34B which has 60 layers. This requires 120 calls to cuMemMap + cuMemSetAccess each of which takes about 40 microseconds. Therefore, growing the KV-cache of one request by new pages (at all layers) adds about 5 millisecond latency to the corresponding iteration. Further, the latency overhead grows proportional to the number of requests that need new pages in a given iteration. We propose the following optimizations to hide the latency of allocation:
Overlapping memory allocation with compute
@@
We leverage the predictability of memory demand to overlap memory allocation with computation. In particular, note that each iteration produces a single output token for every decode request. Therefore, memory demand for a decode iteration is known ahead-of-time. Further, in the decode phase, a request requires at most one new page. vAttention keeps track of the current context length and how many physical memory pages are already mapped for each request. Using this information, it determines when a request would need a new page and uses a background thread to allocate a new page when the preceding iteration is executing. For example, consider that a request R1 would require a new page in iteration i. When the serving framework invokes step API in iteration i-1, vAttention launches a background thread that maps physical memory pages for iteration i. Since iteration latency is typically in the range of 10s-100s of milliseconds, the background thread has enough time to prepare physical memory mappings for an iteration before it starts executing. This way, vAttention hides the latency of CUDA APIs by mapping physical pages in the KV-cache tensors out of the critical path. Note that in every iteration, step API still needs to ensure that physical pages required for the current iteration are actually mapped. If not, required pages are mapped synchronously.
Deferred reclamation + eager allocation
@@
We observe that allocating physical memory for a new request can be avoided in many cases. Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1. To avoid allocating new pages to R2 from scratch, vAttention simply defers the reclamation of R1’s pages and assigns R1’s reqId to R2. This way, R2 uses the same tensors for its KV-cache that R1 was using – which are already backed by physical pages (for security reasons, we still zero-fill such pages before allowing a new request to use them). Therefore, new pages are required only if the context length of R2 is bigger than that of R1.
@@
We further optimize memory allocation by proactively allocating a small number of physical pages ahead of time. For this purpose, we try to keep a certain number of pages (configurable based on expected workload) mapped into the virtual tensors of one of the inactive reqId. When a new request arrives, we allocate this reqId and identify a new reqId to be allocated next and map physical pages for it. In most cases, these eager allocations obviate the need to allocate physical pages in the critical path of prefill execution. Finally, we trigger memory reclamation only when the number of physical pages cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory). We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
@@ 6. Evaluation
Figure 6. Prefill throughput. vAttention backed systems outperform the paged counterparts of both FlashAttention and FlashInfer. Throughput for longer contexts is lower due to the quadratic complexity of prefill attention.
Our evaluation seeks to answer the following questions:
  • How does vAttention perform for prefill and decode phases (§6.1, §6.2)?
  • How vAttention impacts LLM serving throughput (§6.3)?
  • What is the effect of each of our optimizations (§6.4)?
Models and hardware: We evaluate three models Yi-6B (yi-6b-200k-hf, ), Llama-3-8B (llama-8b-hf, ) and Yi-34B (yi-34b-200k-hf, ), using a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 4). Each GPU has 80GB physical memory. We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B.
Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different (sarathiserve2024, ; splitfuse2024, ; distserve2024, ). Attention kernels used for these two phases are also different and hence we also evaluate them separately. The prefill phase requires one time memory allocation potentially spanning multiple pages. In comparison, the decode phase requires incremental memory allocation over the lifetime of a request (vllmsosp, ). We define throughput as the number of tokens processed (or generated) per second.
Serving framework: For a fair comparison, we use vLLM v0.2.7 as a common serving framework in all our experiments. We integrated state-of-the-art kernel libraries of both FlashAttention v2.5.9 (flashattention, ; flashattention2, ; fagithub, ) and FlashInfer v0.4.0 (figithub, ) as attention back-ends into vLLM, and further added support for dynamic memory allocation via vAttention to their non-paged kernels. While FlashAttention and FlashInfer are both based on the same underlying techniques (e.g., FlashDecoding (flashdecoding, )), they use different Block-Table formats; the former use a simple lookup table whereas the latter uses a compressed Block-Table to aid fetching KV-cache block locations.
Baselines: We compare performance obtained by using the non-paged attention kernels (backed by vAttention for dynamic memory allocation), and their paged counterparts. In addition, we also compare against vLLM’s decode kernel (note that vLLM does not have a paged prefill kernel). For a fair comparison, we also profiled each system to find its best performing configuration. Accordingly, we set KV-cache block size to 16 for both vLLM and FlashInfer, and 256 for FlashAttention. Using a higher block size for vLLM increases its kernel latency by up to 3×3$\times$3 × as shown in Figure 3, and using a smaller block size for FlashAttention paged kernel increases its latency by up to $9\%$ (Table 9). We find that the choice of page size does not affect vAttention as long as fragmentation is not a concern (Figure 11).
Model Context Length FA_Paged FA_vAttention FI_Paged FI_vAttention
@@
Figure 7. Decode throughput. FA_vAttention is on par with FA_Paged (note the overlapping lines) which is the best among all PagedAttention based alternatives, while outperforming FI_Paged and vLLM.
@@ 6.1. Prefill Evaluation
We evaluate four configs for prefill: FA_Paged, FI_Paged, FA_vAttention and FI_vAttention. Configurations with the “_Paged” suffix represent the PagedAttention-based kernel and those with “_vAttention” use the non-paged kernels of the respective library. Figure 6 shows the prefill throughput for Yi-6B, Llama-3-8B and Yi-34B. We summarize our key findings below.
@@
Small contexts: For small contexts, prefill cost is dominated by the linear operators i.e., attention’s contribution is relatively low (sarathi2023, ). Hence, even though vAttention helps speed up attention computation, the throughput of both paged and vAttention back-ends is nearly identical in FlashAttention. However, for FlashInfer, we find that using vAttention helps improve prefill throughput even for small contexts. This is because FlashInfer incurs various other sources of overhead in the paged version. First, appending a new K or V tensor to the KV-cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time (the copy operation has been optimized for FlashAttention by vLLM (vllm-copy-kernels, )). Second, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration. vAttention avoids such overheads because it maintains KV-cache’s virtual contiguity, eliminating the need for a Block-Table.
@@
Long contexts: The contribution of attention computation becomes significant at 16K and higher context lengths in our experiments. Therefore, even for FlashAttention back-end, vAttention outperforms the paged counterpart. For example, at context length 192K, FA_vAttention outperforms FA_Paged by 1.24×1.24$\times$1.24 ×, 1.26×1.26$\times$1.26 × and 1.24×1.24$\times$1.24 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. Similarly, for FlashInfer back-ends, FI_vAttention improves prefill throughput by up to 1.25×1.25$\times$1.25 × and 1.36×1.36$\times$1.36 × for Yi-6B and Llama-3-8B (context length 16K), and 1.17×1.17$\times$1.17 × for Yi-34B (context length 32K).
@@
Attention time: vAttention’s improvement in prefill throughput is primarily due to faster attention kernels enabled by a (virtually) contiguous KV-cache. This is because the prefill phase of a long prompt is primarily dominated by attention computation, as can be observed by comparing the numbers inside parenthesis with the total prefill completion time in Table 5. For FlashAttention, nearly all the gains of vAttention are due to faster attention kernels e.g., prefill gains of 1.5 seconds, 7.4 seconds and 16.9 seconds for Yi-6B are almost entirely due to gains in attention computation. vAttention enabled kernels also help with the FlashInfer back-end. In addition, FI_Paged also has other sources of overheads e.g., in the 14 seconds of total savings (Yi-34B, 192K context), only 7 seconds is due to attention and the rest is due to other sources.
Figure 8. Makespan of serving a set of 50 requests with varying context lengths and the ratio of prefill to decode tokens.
6.2. Decode Evaluation
For decodes, in addition to FA_Paged and FI_Paged, we also evaluate the throughput obtained with vLLM’s decode kernel (the first ever kernel to support PagedAttention). For vAttention, we use FlashAttention’s non-paged kernel. Unfortunately FlashInfer’s non-paged decode kernel has significantly higher latency compared to all the other kernels we used in these experiments (up to 14.6×14.6$\times$14.6 ×, Table 10). Hence, while vAttention supports dynamic memory allocation for FlashInfer’s non-paged decode kernel, we omit it for evaluation in this section (note that vAttention is responsible only for allocating memory – it cannot make a slow kernel fast).
Figure 7 shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch size up to 32 (except for Yi-34B which runs out of memory for batch size 32). We set the initial context length of each request to 16K tokens and calculate decode throughput based on the mean latency of 400 decode iterations (see Figure 7).
First, vAttention is on par with the best of PagedAttention as shown by FA_Paged and FA_vAttention in Figure 7. In comparison, FI_Paged has somewhat lower throughput and vLLM is the worst for all models and configurations. For example, FA_Paged and FA_vAttention outperform vLLM by up to 1.99×1.$9$, 1.58×1.58$\times$1.58 × and 1.53×1.53$\times$1.53 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. The primary reason is that vLLM’s decode kernel has significantly higher latency than the FlashAttention based kernels; while FlashAttention has continuously adopted new optimizations (e.g., FlashDecoding (flashdecoding, )), vLLM has lagged behind. For example, Table 6 shows that vLLM’s PagedAttention kernel incurs up to 2.8×2.8$\times$2.8 ×, 1.5×1.5$\times$1.5 ×, and 2.5×2.5$\times$2.5 × higher latency than FlashAttention kernels for Yi-6B, Llama-3-8B and Yi-34B. This is despite vLLM being in an actively maintained open-source serving stack, as well as being used by various companies for serving LLMs. This is an important result that underlines the importance of adopting new optimizations.
Model BS vLLM FA_Paged FI_Paged FA_vAttention

### v3

@@ 5.3.1. Initial setup
4:vattention.init(config_params)
@@
5.3.2. Scheduling a new request
When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid (line 8). All subsequent memory management operations of the request are tagged with this reqId.
5.3.3. Model execution
Before scheduling a batch for execution, the framework needs to ensure that the KV cache sub-tensors of each active request are backed by physical memory (Figure 5(b) and (c)). For this purpose, before dispatching the first kernel of an iteration to the GPU, the framework invokes the step API (line 13), specifying the current context length of each request (context length is set to 0 for each inactive reqId). Internally, vAttention ensures that enough physical pages are mapped for each active reqId before returning execution back to the framework. If vAttention cannot satisfy the memory demand, it returns with a failure in response to which a serving framework can preempt one or more requests to allow forward progress (this is similar to vLLM’s default behavior). We leave more sophisticated policies such as swapping out KV cache to CPU memory as future work.
Depending on whether a request is in the prefill phase or decode phase, different amount of physical memory may need to be mapped for a given iteration. The prefill phase processes the input tokens of a given prompt in parallel. Therefore, the amount of physical memory needed to be mapped depends on the number of prompt tokens being scheduled. If the total K cache size of all prompt tokens at one layer of the model is $s$ and page-group size is $t$, then each worker needs to ensure that at least $(s+t-1)/t$ page-groups are mapped in each of the $2\times N$ KV cache sub-tensors of the given reqId.
For a request in the decode phase, the number of new page-groups required is at most one per virtual tensor. This is because each iteration produces only one output token for a request. vAttention internally tracks the number of page-groups mapped for each request and maps new page-groups only when prior page-groups are about to be exhausted.
5.3.4. Request completion
A request terminates when it reaches user specified or the maximum context length supported by the model, or when the model produces a special end-of-sequence token. The framework notifies vAttention of a request’s completion with free_reqid (line 19). Internally, vAttention may unmap the physical pages of a completed request or defer them to be freed later (§6.1.2).
Supporting continuous batching: Continuous batching poses one challenge in computing attention in our design. When a request from somewhere in the middle of a batch exits, it creates an unused hole in the virtual tensors of KV cache. This layout is not supported by implementations that expect the query (Q) and KV cache to be of the same size in batch ($B$) dimension. However, FlashAttention provides rich API support to address this issue; its argument cache_batch_idx allows Q and KV cache to have different batch sizes and also be arranged in arbitrary order (i.e., batch index 0 in Q can be mapped to batch index 1 in KV cache). vAttention benefits from this API support in terms of both performance and ease of programming; when the request composition of a batch changes, we update cache_batch_idx of running requests such that their Q tensors map to their respective KV cache based on their reqId.
6. Optimizations
@@
There are two challenges in using CUDA virtual memory support for serving LLMs. First, invoking CUDA VMM APIs at runtime incurs high latency. Second, cuMemCreate currently allocates memory only at the granularity of large pages, i.e., multiples of 2MB. Use of large pages can waste physical memory due to internal fragmentation. This section details a set of simple-yet-effective optimizations that we introduce to overcome these challenges.
6.1. Hiding Latency of Memory Allocation
@@
The serving framework invokes the step API in every iteration. The latency of step depends on the number of page-groups that need to be mapped into KV cache. Consider, for example, that the KV cache of one request needs to be extended for Yi-34B which has 60 layers. This requires 120 calls to cuMemMap + cuMemSetAccess each of which takes about 40 microseconds. Therefore, growing the KV cache of one request by new page-groups (two per layer) adds about 5 millisecond latency to the corresponding iteration. The latency overhead grows proportional to the number of requests that need new page-groups in a given iteration. We propose the following optimizations to hide this latency:
6.1.1. Overlapping memory allocation with compute (decode phase)
@@
We leverage the predictability of memory demand to overlap memory allocation with computation. In particular, note that each iteration produces a single output token for every decode request. Therefore, memory demand for a decode iteration is known ahead-of-time. Further, in the decode phase, a request requires at most one new page-group for each of its virtual tensors. vAttention keeps track of the current context length and how much physical memory is already mapped for each request. Using this information, it determines when a request would need more memory and uses a background thread to allocate new page-groups when the preceding iteration is executing. For example, consider that a request R1 would require more physical memory in iteration i. When the serving framework invokes step API in iteration i-1, vAttention launches a background thread that maps page-groups for iteration i. Since per-iteration latency is typically in the range of 10s-100s of milliseconds, the background thread has enough time to prepare physical memory mappings for an iteration before it starts executing. This way, vAttention hides the latency of CUDA APIs by performing memory allocations out of the critical path.
6.1.2. Deferred reclamation + eager allocation (prefill phase)
@@
We observe that allocating physical memory for the prefill phase can be avoided in many cases. Consider that a request R1 completed in iteration i and a new request R2 joins the running batch in iteration i+1. To avoid allocating page-groups to R2 from scratch, vAttention simply defers the reclamation of R1’s page-groups (Figure 5(d)) and assigns R1’s reqId to R2. This way, R2 uses the same tensors for its KV cache that R1 was using – which are already backed by physical pages (Figure 5(e)). Therefore, new allocations are required only if R2’s context length is higher than R1.
@@
We further optimize the prefill phase by proactively allocating a small number of page-groups ahead of time. For this purpose, we try to keep a certain number of page-groups mapped into the virtual tensors of one of the inactive reqId. When a new request arrives, we allocate this reqId. At the same time, we identify a new reqId to be allocated next and eagerly map physical page-groups for it. In most cases, these eager allocations obviate the need to allocate physical memory in the critical path of prefill execution. Finally, we trigger memory reclamation only when the number of page-groups cached in vAttention falls below a certain threshold (e.g., less than 10% of GPU memory). We delegate both deferred reclamation and eager allocation to the background thread that the step API spawns.
Figure 7. Prefill throughput. vAttention backed systems outperform the paged counterparts of both FlashAttention-2 and FlashInfer. Throughput for longer contexts is lower due to the quadratic complexity of prefill attention.
6.2. Mitigating Internal Fragmentation
We mitigate internal fragmentation by reducing the granularity of physical memory allocation. NVIDIA GPUs natively support at least three page sizes: 4KB, 64KB and 2MB (Pratheek et al., 2022; Nayak et al., 2021; Zhang et al., 2023a; pas, 2024). Therefore, in principal, physical memory can be allocated in any multiple of 4KB sizes. The simplest way to achieve this would be to extend the existing CUDA VMM APIs (listed in Table 3) to also support allocating smaller pages (similar to how mmap in Linux supports multiple page sizes (Kwon et al., 2016; Panwar et al., 2019)). Unfortunately, the CUDA VMM APIs are implemented in the closed-source NVIDIA drivers which makes it impossible for us to modify their implementation.
Fortunately, some part of NVIDIA drivers (particularly related to unified memory management) is open-source. Therefore, we implement a new set of APIs in the open-source NVIDIA drivers to mimic the same functionality that existing CUDA APIs provide but with support for multiple page sizes. The second column in Table 3 shows our new APIs: most of our APIs have a one-to-one relationship with existing CUDA APIs except for vMemMap that combines the functionality of cuMemMap and cuMemSetAccess, and vMemRelease that combines the functionality of cuMemUnmap and cuMemRelease for simplicity. In contrast to CUDA VMM APIs, our APIs allocate physical memory in 64KB, 128KB and 256KB sized page-groups. A serving framework can configure a desired page-group size in vAttention while initializing it (we use the standard 2MB pages if the configured page-group size is 2MB). Table 3 shows the latency of each API with different page-group sizes.
@@ 7. Evaluation
Our evaluation answers the following questions:
  • How does vAttention impact the performance of attention kernels and the overall performance of prefill and decode phases (§7.1, §7.2)?
  • How does vAttention impact the end-to-end LLM serving throughput (§7.3, §7.4, §7.5)?
  • What is the effect of each of our optimizations (§7.6)?
Models and hardware: We evaluate three models Yi-6B (yi-, 2024b), Llama-3-8B (lla, 2024) and Yi-34B (yi-, 2024a). We conduct most of our evaluation on a single NVIDIA A100 GPU for Yi-6B, and two NVLink-connected A100 GPUs for Llama-3-8B and Yi-34B (see Table 5). Each GPU has 80GB physical memory. We use tensor-parallelism degree of two (TP-2) for both Llama-3-8B and Yi-34B. To demonstrate the portability benefit of vAttention, we use 1–2 H100 GPUs.
Evaluation methodology: The computation and memory allocation pattern of the prefill and decode phases is substantially different (Agrawal et al., 2024b; Holmes et al., 2024; Zhong et al., 2024). Attention kernels used for these two phases are also different and hence we evaluate them separately. The prefill phase requires one time memory allocation potentially spanning multiple pages. In comparison, the decode phase requires incremental memory allocation over the lifetime of a request (Kwon et al., 2023). We define prefill throughput as the number of prompt tokens processed per second, and decode throughput as the number of tokens generated per second.
@@
Serving framework: For a fair comparison, we use vLLM v0.2.7 as a common serving framework in all our experiments. We integrated state-of-the-art kernel libraries of both FlashAttention-2 v2.5.9 (Dao et al., 2024; Dao, 2023; fag, 2022) and FlashInfer v0.4.0 (fig, 2023) as attention back-ends into vLLM, and further added support for dynamic memory allocation via vAttention to their non-paged kernels. While FlashAttention-2 and FlashInfer are both based on the same underlying techniques (e.g., FlashDecoding (fla, 2023)), they use different Block-Table formats; the former use a simple lookup table whereas the latter uses a compressed Block-Table to optimize lookups.
Baselines: We compare performance obtained by using the non-paged attention kernels (backed by vAttention for dynamic memory allocation), and their paged counterparts. In addition, we also compare against vLLM’s decode kernel (note that vLLM does not have a paged prefill kernel). We also profiled each system to find its best performing configuration. Accordingly, we set KV cache block size to 16 for both vLLM and FlashInfer, and 256 for FlashAttention-2. Using a higher block size for vLLM increases its kernel latency by up to 1.9×1.9$\times$1.9 × as shown in Figure 3, and using a smaller block size for FlashAttention-2 paged kernel increases its latency by up to $9\%$. We find that the choice of page size does not affect vAttention as long as fragmentation is not a concern.
Model Context Length FlashAttention-2 FlashInfer
Paged vAttention Paged vAttention
@@ 7.1. Prefill Evaluation
We evaluate 4 configurations for prefill: FA2_Paged, FI_Paged, FA2_vAttention and FI_vAttention. Configurations with the “_Paged” suffix represent the PagedAttention-based kernel and those with “_vAttention” use the non-paged kernels of the respective library. Figure 7 shows the prefill throughput for our models. We summarize key findings below.
@@
Small contexts: For small contexts, prefill cost is dominated by the linear operators, i.e., attention’s contribution is relatively low (Agrawal et al., 2023). Hence, even though vAttention helps speed up attention computation, the throughput of both paged and vAttention back-ends is nearly identical in FlashAttention-2. However, for FlashInfer, we find that using vAttention helps improve prefill throughput even for small contexts. This is because FlashInfer incurs various other sources of overhead in the paged version. First, appending a new K or V tensor to the KV cache requires a single tensor copy operation in vAttention, whereas in a paged implementation, it requires appending one block at a time (the copy operation has been optimized for FlashAttention-2 by vLLM (vll, 2024a)). Second, FlashInfer involves creation and deletion of a few objects for its compressed Block-Tables in every iteration. vAttention avoids such overheads because it maintains KV cache’s virtual contiguity, eliminating the need for a Block-Table.
Figure 8. Decode throughput. FA2_vAttention is on par with FA2_Paged (note the overlapping lines) which is the best among all PagedAttention based alternatives, while outperforming FI_Paged and vLLM.
@@
Long contexts: The contribution of attention computation becomes significant at 16K and higher context lengths in our experiments. Therefore, even for FlashAttention-2 back-end, vAttention outperforms the paged counterpart. For example, at context length 192K, FA2_vAttention outperforms FA2_Paged by 1.24×1.24$\times$1.24 ×, 1.26×1.26$\times$1.26 × and 1.24×1.24$\times$1.24 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. Similarly, for FlashInfer back-ends, FI_vAttention improves prefill throughput by up to 1.25×1.25$\times$1.25 × and 1.36×1.36$\times$1.36 × for Yi-6B and Llama-3-8B (context length 16K), and 1.17×1.17$\times$1.17 × for Yi-34B (context length 32K).
@@
Attention time: vAttention’s improvement in prefill throughput is primarily due to faster attention kernels enabled by a (virtually) contiguous KV cache. This is because the prefill phase of a long prompt is primarily dominated by attention computation, as can be observed by comparing the numbers inside parenthesis with total prefill completion time in Table 6. For FlashAttention-2, nearly all the gains of vAttention are due to faster attention kernels, e.g., prefill gains of 1.5 seconds, 7.4 seconds and 16.9 seconds for Yi-6B are all due to gains in attention computation. vAttention enabled kernels also help with the FlashInfer back-end. In addition, FI_Paged also has other sources of overheads, e.g., in the 14 seconds of total savings (Yi-34B, 192K context), only 7 seconds is due to attention and the rest is due to other sources.
Model BS vLLM FA2_Paged FI_Paged FA2_vAttention
@@
7.2. Decode Evaluation
For decodes, in addition to FA2_Paged and FI_Paged, we also evaluate the throughput obtained with vLLM’s decode kernel (the first ever kernel to support PagedAttention). For vAttention, we use FlashAttention-2’s non-paged kernel. Unfortunately FlashInfer’s non-paged decode kernel has significantly higher latency (up to 14.6×14.6$\times$14.6 ×) compared to all these other kernels. Hence, while vAttention can support dynamic memory allocation for FlashInfer’s non-paged decode kernel, we omit it for evaluation in this section.
Figure 8 shows the decode throughput of Yi-6B, Llama-3-8B and Yi-34B with varying batch size up to 32 (except for Yi-34B which runs out of memory for batch size 32). We set the initial context length of each request to 16K tokens and calculate decode throughput based on the mean latency of 400 decode iterations.
First, vAttention is on par with the best of PagedAttention as shown by FA2_Paged and FA2_vAttention in Figure 8. In comparison, FI_Paged has somewhat lower throughput and vLLM is the worst for all models and configurations. For example, FA2_Paged and FA2_vAttention outperform vLLM by up to 1.99×1.$9$, 1.58×1.58$\times$1.58 × and 1.53×1.53$\times$1.53 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. The primary reason is that vLLM’s decode kernel has significantly higher latency than the FlashAttention-2 based kernels; while FlashAttention-2 has continuously adopted new optimizations (e.g., FlashDecoding (fla, 2023)), vLLM has lagged behind. For example, Table 7 shows that vLLM’s PagedAttention kernel incurs up to 2.8×2.8$\times$2.8 ×, 1.5×1.5$\times$1.5 ×, and 2.5×2.5$\times$2.5 × higher latency than FlashAttention-2 kernels for Yi-6B, Llama-3-8B and Yi-34B. This is despite vLLM being in an actively maintained open-source serving stack, as well as being used by various companies for serving LLMs. This is an important result that underlines the importance of low software complexity and portability.